OpenStudy (anonymous):

Given that h(x) = 3x − 19, find the value of x that makes h(x) = 71 I thinks its A but i'm not sure 30 43 52 194

OpenStudy (alexandervonhumboldt2):

here is now to do it: h(x) = 3x − 19 h(x)=71 71=3x-19 3x=90 x=90/3=30 so you are correct

OpenStudy (anonymous):

thanks

OpenStudy (alexandervonhumboldt2):

you are welcome!